# SPARK: Sim-ready Part-level Articulated Reconstruction with VLM Knowledge

### v1

@@ Abstract
Articulated 3D objects are critical for embodied AI, robotics, and interactive scene understanding, yet creating simulation-ready assets remains labor-intensive and requires expert modeling of part hierarchies and motion structures. We introduce SPARK, a framework for reconstructing physically consistent, kinematic part-level articulated objects from a single RGB image. Given an input image, we first leverage VLMs to extract coarse URDF parameters and generate part-level reference images. We then integrate the part-image guidance and the inferred structure graph into a generative diffusion transformer to synthesize consistent part and complete shapes of articulated objects. To further refine the URDF parameters, we incorporate differentiable forward kinematics and differentiable rendering to optimize joint types, axes, and origins under VLM-generated open-state supervision. Extensive experiments show that SPARK produces high-quality, simulation-ready articulated assets across diverse categories, enabling downstream applications such as robotic manipulation and interaction modeling. Project page: https://heyumeng.com/SPARK/index.html
@@ 3.2 Part-Articulated Object Generation
We begin by extracting visual guidance for each part. A shared DINOv2 encoder [oquab2023dinov2] maps every part image $\mathbf{r}_{k}$ and the replicated global image $I_{0}$ into token sequences, producing local embeddings $\mathbf{E}^{\mathrm{loc}}_{k}\!\in\!\mathbb{R}^{L\times d}$ and a global embedding $\mathbf{E}^{\mathrm{glob}}\!\in\!\mathbb{R}^{L\times d}$. These embeddings provide part-specific and object-level conditioning for the generative process. To integrate these image features into 3D reconstruction, the DiT maintains its own learnable latent tokens that represent the geometry to be generated. Let an object contain $K$ parts, each represented by $N$ latent tokens. We stack them as $Z=[Z_{1};\dots;Z_{K}]\in\mathbb{R}^{NK\times C}$, where $Z_{i}\in\mathbb{R}^{N\times C}$. During denoising, the model fuses the image embeddings with the latent geometry tokens through alternating local and global cross-attention blocks. Local blocks query the corresponding part embedding $\mathbf{E}^{\mathrm{loc}}_{k}$, ensuring each part attends only to its own visual reference, while global blocks query $\mathbf{E}^{\mathrm{glob}}$ to inject whole-object context shared across parts. Within these blocks, latent-to-latent attention is computed using the local map $\mathbf{A}^{\mathrm{local}}_{i}=\mathrm{softmax}\!\big(Z_{i}Z_{i}^{\top}/\sqrt{C}\big)\in\mathbb{R}^{N\times N}$ and the global map $\mathbf{A}^{\mathrm{global}}=\mathrm{softmax}\!\big(ZZ^{\top}/\sqrt{C}\big)\in\mathbb{R}^{NK\times NK}$, where $\mathbf{A}$ denotes the attention weights that propagate information within each part and across the full object. To incorporate structural guidance from the VLM priors, we introduce a hierarchical attention mechanism that exchanges information between each parent–child pair. A parent index map $\pi:\{1,\dots,K\}\!\to\!\{-1,1,\dots,K\}$ defines the link hierarchy; $\pi(k)=-1$ for root. Let $\mathcal{P}(u)$ and $\mathcal{C}(u)$ denote the token sets of the parent and child of token $u$, respectively. Child tokens attend only to parent tokens via
@@ 4.3 Ablation Study
We conduct an ablation study on the key components of our framework: part guidance, data augmentation, and joint optimization, to evaluate their contributions to mesh reconstruction and URDF parameter estimation. VLM-driven part guidance provides strong visual and structural priors. As shown in Fig. 6, removing part guidance leads to missing doors in the cabinet. Without data augmentation, the model overfits to limited states and may confuse visually similar parts (e.g., misidentifying two cabinet doors). Incorporating both VLM-derived guidance and data augmentation yields cleaner part boundaries, more consistent decomposition, and higher-quality reconstructions. As reported in Table 3, these components together achieve the best geometric accuracy.
@@
Joint optimization further refines URDF parameters, improving joint-type prediction, joint-axis orientation, and joint pivot estimation (Table 4). Removing optimization leads to misaligned or drifting moving parts, such as shifted pivot points (Fig. 6).
@@ 5 Conclusion
[…]
Supplementary Material
6 Network Architecture Details
@@
6.1 Latent Representation and Image Encoders
@@
$E^{\text{loc}}_{k}\in\mathbb{R}^{L\times d_{\text{img}}},\hskip 28.80008ptE^{\text{glob}}\in\mathbb{R}^{L\times d_{\text{img}}},$
@@
6.2 Diffusion Transformer Backbone
@@
where $\mathbf{Z}_{t}$ is the noisy latent at time $t$ (defined in Sec. 6.5), $\phi_{\text{time}}$ is a sinusoidal+MLP timestep embedding, and $\cdot$$\cdot\,$] denotes concatenation along the token dimension.
@@
6.3 Multi-Level Attention and Image Conditioning
@@
$\mathbf{A}^{\text{self}}=\mathrm{softmax}\big(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{D}\big),\hskip 28.80008pt\mathbf{H}^{\prime}=\mathbf{H}+\mathbf{A}^{\text{self}}\mathbf{V},$ (7)
@@
$\mathbf{A}^{\text{glob}}=\mathrm{softmax}\big(\mathbf{Q}\mathbf{K}^{\text{glob}\,\top}/\sqrt{D}\big),\hskip 28.80008pt\mathbf{H}^{\prime\prime}=\mathbf{H}^{\prime}+\mathbf{A}^{\text{glob}}\mathbf{V}^{\text{glob}},$ (8)
@@
$\mathbf{A}^{\text{loc}}_{k}=\mathrm{softmax}\big(\mathbf{Q}_{k}\mathbf{K}^{\text{loc}\,\top}_{k}/\sqrt{D}\big),\hskip 28.80008pt\mathbf{H}^{\prime\prime}_{k}=\mathbf{H}^{\prime}_{k}+\mathbf{A}^{\text{loc}}_{k}\mathbf{V}^{\text{loc}}_{k}.$ (9)
@@
6.4 Hierarchy Attention Implementation
@@
6.5 Rectified Flow Training Objective
@@
$\mathbf{Z}_{0}=\begin{bmatrix}\mathbf{z}_{1,0}\\ \vdots\\ \mathbf{z}_{K,0}\end{bmatrix},\hskip 28.80008pt\mathbf{Z}_{1}=\begin{bmatrix}\mathbf{z}_{1,1}\\ \vdots\\ \mathbf{z}_{K,1}\end{bmatrix}$.
@@
6.6 Classifier-Free Guidance
@@
7 Kinematic-Part Mesh Merging
@@
7.1 URDF-Driven Link Grouping
@@
7.2 Per-Link Geometry Cleaning and Merging
@@
    • In both cases, we assign a uniform gray face color [128, 128, 128, 255] to decouple our geometry pipeline from the original textures. Later, we re-render per-part images using our own lighting and camera setup (Sec. 9), so we do not rely on baked-in materials.
@@
This procedure yields one rigid mesh per URDF link, with all subcomponents fused into the same local frame. We intentionally do not perform heavy processing at this stage (e.g., no automatic repairs or decimation) to preserve the original geometry as much as possible; watertight voxelization and cleanup are deferred to the next stage (Sec. 8).
8 Watertight Part Preprocessing
Our network operates on per-part surface samples extracted from mesh geometry (Sec. 6). In practice, raw CAD / reconstruction data often contain small gaps, self-intersections, or open boundaries, which lead to unstable surface sampling and inconsistent volumes. Before training, we therefore convert every part mesh into a watertight surface via voxelization and marching cubes, combined with thin-part guards and aggressive post-cleaning. Unless otherwise stated, all experiments use a target voxel resolution of R=200R{=}200 along the largest object axis.
8.1 Per-part Voxelization and Pitch Selection
@@
$\Delta_{\text{res}}=\frac{\max(\mathbf{e})}{R},\hskip 28.80008pt\Delta_{\text{thin}}=\frac{\min(\mathbf{e})}{N_{\min}},$ (20)
@@
8.2 Closed Occupancy and Marching Cubes
@@
8.3 Axis-wise Rescaling Back to Original Extents
@@
8.4 Robust Mesh Cleanup and Early Exit for Watertight Parts
@@
9 Articulation-Aware Data Augmentation
After constructing watertight per-part meshes and canonical surface samples (Sec. 8), we augment the dataset by rendering each articulated object at additional joint configurations while keeping the underlying geometry and part decomposition fixed. Concretely, for every mobility.urdf with at least one revolute joint, we synthesize two extra variants:
@@ 9 Articulation-Aware Data Augmentation
Each variant reuses the same watertight mesh and part-level surface samples but provides a new RGB rendering at a different articulation state. This encourages the model to become robust to joint motion while still reconstructing a consistent canonical geometry for each object.:contentReference[oaicite:0]index=0
9.1 URDF-Based Multi-Joint Forward Kinematics
@@
$T^{\text{world}}_{\text{vis}}(r)=T^{\text{vis}}_{r},\hskip 28.80008ptT^{\text{world}}_{\text{link}}(r)=I_{4},$
@@
9.2 Sampling Joint Angles: Reference, Mid, and Max Poses
@@
$\theta^{\text{ref}}_{j}=0,\hskip 28.80008pt\theta^{\text{mid}}_{j}=\frac{1}{2}\,\theta^{\max}_{j},\hskip 28.80008pt\theta^{\text{max}}_{j}=\theta^{\max}_{j}.$
@@
9.3 Consistent Normalization and Rendering Settings

### v2

@@ Abstract
Articulated 3D objects are critical for embodied AI, robotics, and interactive scene understanding, yet creating simulation-ready assets remains labor-intensive and requires expert modeling of part hierarchies and motion structures. We introduce SPARK, a framework for reconstructing physically consistent, kinematic part-level articulated objects from a single RGB image. Given an input image, we first leverage VLMs to extract coarse URDF parameters and generate part-level reference images. We then integrate the part-image guidance and the inferred structure graph into a generative diffusion transformer to synthesize consistent part and complete shapes of articulated objects. To further refine the URDF parameters, we incorporate differentiable forward kinematics and differentiable rendering to optimize joint types, axes, and origins under VLM-generated open-state supervision. Extensive experiments show that SPARK produces high-quality, simulation-ready articulated assets across diverse categories, enabling downstream applications such as robotic manipulation and interaction modeling. Project page: https://heyumeng.com/SPARK/index.html.
@@
* equal contribution. 1 UCLA, 2 USC, 3 University of Utah. heyumeng@usc.edu, jiayin_lu,yingjiang,cffjiang@ucla.edu, yin.yang@utah.edu
@@ 3.2 Part-Articulated Object Generation
We begin by extracting visual guidance for each part. A shared DINOv2 encoder [oquab2023dinov2] maps every part image $\mathbf{r}_{k}$ and the replicated global image $I_{0}$ into token sequences, producing local embeddings $\mathbf{E}^{\mathrm{loc}}_{k}\!\in\!\mathbb{R}^{L\times d}$ and a global embedding $\mathbf{E}^{\mathrm{glob}}\!\in\!\mathbb{R}^{L\times d}$. These embeddings provide part-specific and object-level conditioning for the generative process. To integrate these image features into 3D reconstruction, the DiT maintains its own learnable latent tokens that represent the geometry to be generated. Let an object contain $K$ parts, each represented by $N$ latent tokens. We stack them as $Z=[Z_{1};\dots;Z_{K}]\in\mathbb{R}^{NK\times C}$, where $Z_{i}\in\mathbb{R}^{N\times C}$.
During denoising, the model fuses the image embeddings with the latent geometry tokens through alternating local and global cross-attention blocks. Local blocks query the corresponding part embedding $\mathbf{E}^{\mathrm{loc}}_{k}$, ensuring each part attends only to its own visual reference, while global blocks query $\mathbf{E}^{\mathrm{glob}}$ to inject whole-object context shared across parts. Within these blocks, latent-to-latent attention is computed using the local map $\mathbf{A}^{\mathrm{local}}_{i}=\mathrm{softmax}\!\big(Z_{i}Z_{i}^{\top}/\sqrt{C}\big)\in\mathbb{R}^{N\times N}$ and the global map $\mathbf{A}^{\mathrm{global}}=\mathrm{softmax}\!\big(ZZ^{\top}/\sqrt{C}\big)\in\mathbb{R}^{NK\times NK}$, where $\mathbf{A}$ denotes the attention weights that propagate information within each part and across the full object. To incorporate structural guidance from the VLM priors, we introduce a hierarchical attention mechanism that exchanges information between each parent–child pair. A parent index map $\pi:\{1,\dots,K\}\!\to\!\{-1,1,\dots,K\}$ defines the link hierarchy; $\pi(k)=-1$ for root. Let $\mathcal{P}(u)$ and $\mathcal{C}(u)$ denote the token sets of the parent and child of token $u$, respectively. Child tokens attend only to parent tokens via
@@ 4.3 Ablation Study
We conduct an ablation study on the key components of our framework: part guidance, data augmentation, and joint optimization, to evaluate their contributions to mesh reconstruction and URDF parameter estimation. VLM-driven part guidance provides strong visual and structural priors. As shown in Fig. 6, removing part guidance leads to missing doors in the cabinet. Without data augmentation, the model overfits to limited states and may confuse visually similar parts (e.g., misidentifying two cabinet doors). Incorporating both VLM-derived guidance and data augmentation yields cleaner part boundaries, more consistent decomposition, and higher-quality reconstructions. As reported in Table 3, these components together achieve the best geometric accuracy. Joint optimization further refines URDF parameters, improving joint-type prediction, joint-axis orientation, and joint pivot estimation (Table 4). Removing optimization leads to misaligned or drifting moving parts, such as shifted pivot points (Fig. 6).
@@ 5 Conclusion
Appendix
Appendix A Network Architecture Details
@@
A.1 Latent Representation and Image Encoders
@@
$E^{\text{loc}}_{k}\in\mathbb{R}^{L\times d_{\text{img}}},\qquad E^{\text{glob}}\in\mathbb{R}^{L\times d_{\text{img}}},$
@@
A.2 Diffusion Transformer Backbone
@@
where $\mathbf{Z}_{t}$ is the noisy latent at time $t$ (defined in Sec. A.5), $\phi_{\text{time}}$ is a sinusoidal+MLP timestep embedding, and $\cdot$$\cdot\,$] denotes concatenation along the token dimension.
@@
A.3 Multi-Level Attention and Image Conditioning
@@
$\mathbf{A}^{\text{self}}=\mathrm{softmax}\big(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{D}\big),\qquad\mathbf{H}^{\prime}=\mathbf{H}+\mathbf{A}^{\text{self}}\mathbf{V},$ (7)
@@
$\mathbf{A}^{\text{glob}}=\mathrm{softmax}\big(\mathbf{Q}\mathbf{K}^{\text{glob}\,\top}/\sqrt{D}\big),\qquad\mathbf{H}^{\prime\prime}=\mathbf{H}^{\prime}+\mathbf{A}^{\text{glob}}\mathbf{V}^{\text{glob}},$ (8)
@@
$\mathbf{A}^{\text{loc}}_{k}=\mathrm{softmax}\big(\mathbf{Q}_{k}\mathbf{K}^{\text{loc}\,\top}_{k}/\sqrt{D}\big),\qquad\mathbf{H}^{\prime\prime}_{k}=\mathbf{H}^{\prime}_{k}+\mathbf{A}^{\text{loc}}_{k}\mathbf{V}^{\text{loc}}_{k}.$ (9)
@@
A.4 Hierarchy Attention Implementation
@@
A.5 Rectified Flow Training Objective
@@
$\mathbf{Z}_{0}=\begin{bmatrix}\mathbf{z}_{1,0}\\ \vdots\\ \mathbf{z}_{K,0}\end{bmatrix},\qquad\mathbf{Z}_{1}=\begin{bmatrix}\mathbf{z}_{1,1}\\ \vdots\\ \mathbf{z}_{K,1}\end{bmatrix}$.
@@
A.6 Classifier-Free Guidance
@@
Appendix B Kinematic-Part Mesh Merging
@@
B.1 URDF-Driven Link Grouping
@@
B.2 Per-Link Geometry Cleaning and Merging
@@
    • In both cases, we assign a uniform gray face color [128, 128, 128, 255] to decouple our geometry pipeline from the original textures. Later, we re-render per-part images using our own lighting and camera setup (Sec. D), so we do not rely on baked-in materials.
@@
This procedure yields one rigid mesh per URDF link, with all subcomponents fused into the same local frame. We intentionally do not perform heavy processing at this stage (e.g., no automatic repairs or decimation) to preserve the original geometry as much as possible; watertight voxelization and cleanup are deferred to the next stage (Sec. C).
Appendix C Watertight Part Preprocessing
Our network operates on per-part surface samples extracted from mesh geometry (Sec. A). In practice, raw CAD / reconstruction data often contain small gaps, self-intersections, or open boundaries, which lead to unstable surface sampling and inconsistent volumes. Before training, we therefore convert every part mesh into a watertight surface via voxelization and marching cubes, combined with thin-part guards and aggressive post-cleaning. Unless otherwise stated, all experiments use a target voxel resolution of R=200R{=}200 along the largest object axis.
C.1 Per-part Voxelization and Pitch Selection
@@
$\Delta_{\text{res}}=\frac{\max(\mathbf{e})}{R},\qquad\Delta_{\text{thin}}=\frac{\min(\mathbf{e})}{N_{\min}},$ (20)
@@
C.2 Closed Occupancy and Marching Cubes
@@
C.3 Axis-wise Rescaling Back to Original Extents
@@
C.4 Robust Mesh Cleanup and Early Exit for Watertight Parts
@@
Appendix D Articulation-Aware Data Augmentation
After constructing watertight per-part meshes and canonical surface samples (Sec. C), we augment the dataset by rendering each articulated object at additional joint configurations while keeping the underlying geometry and part decomposition fixed. Concretely, for every mobility.urdf with at least one revolute joint, we synthesize two extra variants:
@@ Appendix D Articulation-Aware Data Augmentation
Each variant reuses the same watertight mesh and part-level surface samples but provides a new RGB rendering at a different articulation state. This encourages the model to become robust to joint motion while still reconstructing a consistent canonical geometry for each object.
D.1 URDF-Based Multi-Joint Forward Kinematics
@@
$T^{\text{world}}_{\text{vis}}(r)=T^{\text{vis}}_{r},\qquad T^{\text{world}}_{\text{link}}(r)=I_{4},$
@@
D.2 Sampling Joint Angles: Reference, Mid, and Max Poses
@@
$\theta^{\text{ref}}_{j}=0,\qquad\theta^{\text{mid}}_{j}=\frac{1}{2}\,\theta^{\max}_{j},\qquad\theta^{\text{max}}_{j}=\theta^{\max}_{j}.$
@@
D.3 Consistent Normalization and Rendering Settings